Let C be square matrix.

i) Check if S = C + CT is symmetric

ii) Check if N + C - CTis skew symmetric

iii) Prove that every square matrix can be written as a sum of skew symmetric matrix and symmetric matrix
Answer #1


sel: (i) s= (400) T-c46c0) =c+c=5. Hotel besoises 3 Sis symmelic matrix, cil -c-ct. NTE (C-CJ) T- C- C =-(CT) --N so 17-N 7